solve the systm of linear equation by substitution
3x+2y=6
y=2x+10

Answer by stanbon(75887)

solve the systm of linear equation by substitution
3x+2y=6
y=2x+10
--------------------
Substitute for "y" and solve for "x":
3x + 2(2x+10) = 6
3x + 4x + 20 = 6
7x = -14
x = -2
--------
Substitute into y = 2x+10 to solve for "y":
y = 2(-2)+10
---
y = 6
------
Solution: (-2,6)
=================
Cheers,
Stan H.
=================